Each of the following samples was obtained from a population with µ = 100 and σ = 10. Find
the z-score corresponding to each sample mean.
a. M = 95 for a sample of n = 4
b. M = 104 for a sample of n = 25
c. M = 103 for a sample of n = 100
1
Expert's answer
2019-05-01T11:09:57-0400

If X has a distribution with mean "\mu," and standard deviation "\sigma," and is approximately normally distributed or "n" is large, then the sample mean M is approximately normally distributed with mean "\mu" and standard error "\sigma/\sqrt{n}."


"z={M-\mu \over \sigma/\sqrt{n}}"

Given that "\mu=100" and "\sigma=10."

a. M = 95 for a sample of n = 4 


"z={95-100 \over 10/\sqrt{4}}=-1"

b. M = 104 for a sample of n = 25 


"z={104-100 \over 10/\sqrt{25}}=2"

c. M = 103 for a sample of n = 100


"z={103-100 \over 10/\sqrt{100}}=3"